# The Machine Learning Control Method for Counterfactual $*{}^{*}$

### v1

@@ Appendix A Bootstrap inference for ATE and CATE confidence intervals
Appendix B Overview of the ML models employed in the application
This appendix provides an overview of the ML algorithms employed in both the simulation study of section 3.3 and the empirical application of section 4. We remark that modeling potential outcomes in the absence of the treatment according to Equation (2) reflects our uncertainty on $Y_{i,t}(0)$ rather than an assumption on the distribution of potential outcomes in the superpopulation, as we already observe the entire population under study. As discussed in the main text, we tested several ML methods to learn the function $f(\cdot)$. Thus, we provide some examples of what the causal effect estimators look like for some of the supervised ML methods that can be employed by the MLCM. Specifically, we show the cases of the four ML algorithms employed in the empirical application.
  • LASSO
    Assuming a linear functional form for $f(\cdot)$ in Equation (2), an estimator for the unobserved potential outcome in the absence of the intervention can be written as,
    $\operatorname{Y}_{i,t_{0}+k|t_{0}}(0)$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+k}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+k}]$
    $=\operatorname{E}\left[\sum_{j=1}^{p}\phi_{j}\operatorname{Y}_{i,t_{0}+k-j}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+k}\right]+\sum_{j=0}^{q}\mathbf{X}_{i,t_{0}+k-j}\widehat{\boldsymbol{\beta}}.$
    This can be recursively computed from $\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)$ as follows:
    $\operatorname{Y}_{i,t_{0}+1|t_{0}}(0)$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+1}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+1}]$
    $=\sum_{j=1}^{p}\hat{\phi}_{j}\operatorname{Y}_{i,t_{0}+1-j}+\sum_{j=0}^{q}\mathbf{X}_{i,t_{0}-j+1}\widehat{\boldsymbol{\beta}}$
    $\operatorname{Y}_{i,t_{0}+2|t_{0}}(0)$ $=\phi_{1}\operatorname{E}\left[\operatorname{Y}_{i,t_{0}+1}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+2}\right]+\sum_{j=2}^{p}\hat{\phi}_{j}\operatorname{Y}_{i,t_{0}+2-j}+\sum_{j=0}^{q}\mathbf{X}_{i,t_{0}-j+2}\widehat{\boldsymbol{\beta}}$
    where ($\hat{\phi}_{1},\dots,\hat{\phi}_{p}$, $\widehat{\boldsymbol{\beta}}$) are estimated as in Tibshirani, (1996) by minimizing a penalized version of the residual sum of squares.
  • Partial Least Squares
    $\operatorname{Y}_{i,t_{0}+k|t_{0}}(0)$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+k}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+k}]$
    $=\operatorname{E}\left[\sum_{j=1}^{p}\phi_{j}\operatorname{Y}_{i,t_{0}+k-j}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+k}\right]+\sum_{j=0}^{q}\mathbf{X}_{i,t_{0}+k-j}\widehat{\boldsymbol{\beta}}^{\text{PLS}}$
    $=\sum_{j=1}^{p}\hat{\phi}_{j}^{\text{PLS}}\operatorname{Y}_{i,t_{0}+1-j}+\sum_{j=0}^{q}\mathbf{X}_{i,t_{0}-j+1}\widehat{\boldsymbol{\beta}}^{\text{PLS}}$
    where ($\hat{\phi}_{1}^{\text{PLS}},\dots,\hat{\phi}_{p}^{\text{PLS}}$, $\widehat{\boldsymbol{\beta}}^{\text{PLS}}$) are estimated as in Geladi and Kowalski, (1986).
  • Stochastic Gradient Boosting
    $\operatorname{Y}_{i,t_{0}+k|t_{0}}(0)$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+k}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+k}]$
    $=\sum_{s=1}^{S}\beta_{s}h_{s}\left(\mathcal{Y}^{(p)}_{i,t_{0}+k-1}(0),\mathbf{X}_{i,t_{0}+k},\mathcal{X}^{(q)}_{i,t_{0}+k-1}\right)$
    where, following Friedman, (2001), $h_{s}$ are the so-called “base” learners chosen to be simple functions of the covariates over $S$ samples. We employ stochastic gradient boosting instead of conventional gradient boosting due to its enhanced approximation accuracy and faster execution speed (Friedman,, 2002).
  • Random Forest
    $\operatorname{Y}_{i,t_{0}+k|t_{0}}(0)$ $=\operatorname{E}[\operatorname{Y}_{i,t_{0}+k}(0)|\mathcal{I}_{t_{0}},\mathbf{X}_{i,t_{0}+k}]$
    $=\frac{1}{A}\sum_{a=1}^{A}T^{(a)}\left(\mathcal{Y}^{(p)}_{i,t_{0}+k-1}(0),\mathbf{X}_{i,t_{0}+k},\mathcal{X}^{(q)}_{i,t_{0}+k-1};R_{Y},\mu\right).$
    In the latter equation, $T(\cdot)$ denotes a single tree, defined as,
    $T\left(\mathcal{Y}^{(p)}_{i,t_{0}+k-1}(0),\mathbf{X}_{i,t_{0}+k},\mathcal{X}^{(q)}_{i,t_{0}+k-1};R_{Y},\mu\right)=\sum_{l=1}^{L}\mu_{l}\mathbb{I}_{\{(\mathcal{Y},X,\mathcal{X}\in R_{l})_{t_{0}+k}\}}$
    where $R_{Y}$ is a binary recursive partition of the covariates space, $l$ is the number of leaf nodes, $\mu_{l}$ are the parameters and $\mathbb{I}_{\{(\mathcal{Y},X,\mathcal{X})_{t_{0}+k}\in R_{l}\}}$ is an indicator function taking value $1$ when unit $i$ (as defined by its covariate set) at time $t_{0}+k$ falls in the partition $R_{l}$. Finally, $A$ is the number of trees forming the forest.
Appendix C Causal estimands in settings with interference between units
Compared to Case (i), in the scenario with violations of the no-interference assumption, the causal estimands are different. In this section, we define causal estimands for the setup in which only a subgroup of units gets treated, but there are likely spillover and general equilibrium effects that prevent using the set of untreated units to build a valid control group (Case (ii) defined above). We can first define the ATT across the treated units at a given point in time.
Definition 6
Definition 7
@@ Appendix D Additional simulation results
In this Section of the Appendix we report the results of all the simulation scenarios that we generated by combining the following parameters: i) number of units in the panel ($N=400$, $N=200$); ii) autoregressive coefficient ($\phi=0.8$, $\phi=1.2)$; iii) standard deviation of $u_{i}$ regulating the amount of heterogeneity between units ($\sigma_{u}=1$, $\sigma_{u}=0.1$).
Table 4 shows the results of an “explosive” simulation scenario corresponding to the one reported in the main text ($N=400$, $\sigma_{u}=1$), but with $\phi=1.2$. Tables 5 and 6 report the results for the remaining combinations with $N=400$ and $\sigma_{u}=0.1$.
Tables 710 show the results of the simulations when the number of units in the panel reduces to $N=200$.
Table 4: Simulation results for $N=400$, $\phi=1.2$ and $\sigma_{u}=1$
@@
Table 5: Simulation results for $N=400$, $\phi=0.8$ and $\sigma_{u}=0.1$
@@
Table 6: Simulation results for $N=400$, $\phi=1.2$ and $\sigma_{u}=0.1$
@@
Table 7: Simulation results for $N=200$ units, $\phi=0.8$ and $\sigma_{u}=1$
@@
Table 8: Simulation results for $N=200$ units, $\phi=0.8$ and $\sigma_{u}=0.1$
@@
Table 9: Simulation results for $N=200$ units, $\phi=1.2$ and $\sigma_{u}=1$
@@
Table 10: Simulation results for $N=200$ units, $\phi=1.2$ and $\sigma_{u}=0.1$
@@ Appendix E Additional application material
Table 11: Definition of the variables included in the initial dataset
@@
Standardized math test score
Standardized math test score of fifth-grade students.
School year 2012/2013-school year 2020/2021
National Institute for the Evaluation of Educational Instruction and Training (INVALSI)
@@
Economic classification dummies
Without specialization, non-manufacturing (touristic), non-manufacturing (non-touristic), made in Italy, other manufacturing
2011
Italian National Institute of Statistics (Istat)
Dummy district
LLM with at least one industrial district
2011
Istat
Regional dummies
A dummy for each of the 20 Italian regions
Istat
Population dummies
$\leq$ 10,000; (10,000; 50,000]; (50,000; 100,000]; (100,000; 500,000]; >500,000
2011
Istat
Share of individuals with a university degree
Share of individuals aged 30-34 with a university degree
2014
Istat
Share of individuals with a high-school degree
Share of individuals aged 30-34 with a high-school degree
2014
Istat
Share of urban area
Urban surface / Total surface
2012
Istat
Share of population in the periphery
Share of population located in municipalities considered as peripheral or ultra-peripheral according to the SNAI classification
2011
Istat
Number of public libraries
Number of public libraries
2016
Istat
Number of visitors to museums
Number of visitors to museums, galleries, archaeological sites, and monuments per 100,000 inhabitants
2015
Istat
Average maximum temperature
30-year average (1980-2010) of maximum air temperature (${}^{\circ}$)
1980-2010
CAMS European Air Quality Reanalysis Data
Average minimum temperature
30-year average (1980-2010) of minimum air temperature (${}^{\circ}$)
1980-2010
CAMS European Air Quality Reanalysis Data
Average total yearly precipitation
30-year average (1980-2010) of total precipitation (mm)
1980-2010
CAMS European Air Quality Reanalysis Data
Turnout referendum
Turnout at the constitutional referendum that was held in Italy on 4 December 2016
2016
Ministry of the Interior
Share of Yes at the referendum
Share of Yes at the constitutional referendum that was held in Italy on 4 December 2016
2016
Ministry of the Interior
Percentage of children aged 0-2 years who have utilized childcare services
Children aged 0-2 years who have utilized childcare services provided by Municipalities (nurseries, micro-nurseries, or integrated and innovative services) / Resident children aged 0-2 years * 100.
2016
Istat
@@
Standardized Italian test score
Standardized Italian test score of fifth-grade students.
School year 2012/2013-school year 2018/2019
INVALSI
Percentage of participants to the math test
Percentage of participation in the math test (pupils who took the test out of expected pupils)
School year 2012/2013-school year 2018/2019
INVALSI
Percentage of participants to the Italian test
Percentage of participation in the Italian test (pupils who took the test out of expected pupils)
School year 2012/2013-school year 2018/2019
INVALSI
Average percentage score for math
Average percentage score for math - adjusted for cheating
School year 2012/2013-school year 2018/2019
INVALSI
Average percentage score for Italian
Average percentage score for Italian - adjusted for cheating
School year 2012/2013-school year 2018/2019
INVALSI
Standard deviation for math
Standard deviation of the percentage score for math
School year 2012/2013-school year 2018/2019
INVALSI
Standard deviation for Italian
Standard deviation of the percentage score for Italian
School year 2012/2013-school year 2018/2019
INVALSI
Log of the Gini index
Log of the Gini index
2013-2019
Ministry of Economy and Finance (MEF)
Share of foreign population
Foreigners / population
2013-2019
Istat
Unemployment rate
Resident population aged 15+ not in employment but currently available for work / Labor force
2013-2019
Istat
Share of graduate mayors
Share of municipalities with a mayor with a university degree
2013-2019
Ministry of the Interior
Share of recycled waste
Share of recycled waste
2013-2019
Italian National Institute for Environmental Protection and Research (Ispra)
Share of workers in manufacturing
Share of workers in manufacturing
2013-2019
Istat
Share of old population
Share of population aged >=65
2013-2019
Istat
Declared income per capita
Declared income per capita
2013-2019
MEF
Log of declared income (total)
Log of declared income (total)
2013-2019
MEF
Share of income for pensions
Share of overall declared income for pensions
2013-2019
MEF
Share of income for employment
Share of overall declared income for employment
2013-2019
MEF
Population
Resident population
2013-2019
Istat
Average price per square meter
Average price per square meter - house
2013-2019
Osservatorio del Mercato Immobiliare – Agenzia delle Entrate
Notes. In our application, time-invariant covariates are either features of the LLM not subject to time changes (e.g. area of the LLM) or features taken at a time before 2017. Data from the 2020/2021 school year represent the first available post-pandemic data point for the standardized math test score, as there was no test for the school year 2019/2020 due to the national lockdown. Following Carlana et al., (2023), math test scores have been standardized for the entire dataset using the 2013-2019 values as a benchmark, with a mean of 0 and a standard deviation of 1.
Table 12: Performances with different numbers of variables
@@
Table 13: Definition of the 20 variables included in the estimation process (random forest)
@@
Notes. These are the 20 variables selected via the panel CV procedure by the preliminary random forest.
Table 14: Definition of the variables included in the data-driven CATEs analysis
@@
COVID-19 impact on the standardized math test score (dependent variable)
Treatment effect of the COVID-19 crisis on the standardized math test score (SD)
School year 2020/2021
Estimated via the MLCM with random forest
Share of income accruing to low-income earners
Share of income accruing to individuals with an overall income $\leq$ 26,000 Euro
2019
MEF
Average download speed of the internet
Average download speed of the internet
2018
AGCOM
Share of households without internet access
Share of households without internet access
2018
AGCOM
Tourism LLM
Dummy variable equal to 1 for LLMs specialized in tourism
2011
Istat
Share of individuals with a university degree
Share of individuals aged 30-34 with a university degree
2015
Istat
Population
Resident population
2019
Istat
Per capita expenditure on gambling
Per capita expenditure on gambling
2019
Unemployment rate
@@
Resident population aged 15+ not in employment but currently available for work
2019
Istat
Excess mortality estimates
@@
Municipality-level excess mortality estimated by applying ML techniques to all-cause deaths data, aggregated at the LLM level
From Feb 21, 2020 to Sep 30, 2020
(Cerqua et al.,, 2021)
Share of temporary contracts
@@
Number of employees with temporary contracts in October divided by the number of employees in October
2015
Istat
Share of jobs in suspended economic activities
@@
Share of jobs in activities suspended in March 2020 by the Italian Government due to the spread of the pandemic
2017
Istat
Per capita income
The amount of money earned per person
2019
MEF
Share of firms having employees in short-time work subsidies
@@
The number of firms with employees receiving short-time work subsidies divided by the universe of firms registered in the Business Register
Average (2015-2018)
Ministry of Labor and Social Policies
Share of population living in peripheral areas
@@
Share of population living in areas defined by Istat as peripheral or ultra-peripheral
Jan 1, 2020
Istat
Dependency ratio
@@
The ratio of those typically not in the labor force (the dependent part, ages 0 to 14 and 65+) and those typically in the labor force (the productive part, ages 15 to 64)
Jan 1, 2020
Istat
Gini index
Gini index
2019
MEF
Average price per square meter
Average price per square meter - house
2019
@@
Osservatorio del Mercato Immobiliare – Agenzia delle Entrate
Figure 6: Unit-level panel placebo test by year
Figure 7: Variable importance ranking – Preliminary random forest
/body>

### v2

@@ Appendix D Additional simulation results
In this Section of the Supplemental Appendix, we report the results of all the simulation scenarios that we generated by combining the following parameters: i) number of units in the panel ($N=400$, $N=200$); ii) autoregressive coefficient ($\phi=0.8$, $\phi=1.2)$; iii) standard deviation of $u_{i}$ regulating the amount of heterogeneity between units ($\sigma_{u}=1$, $\sigma_{u}=0.1$).
Table D.1 shows the results of an “explosive” simulation scenario corresponding to the one reported in the main text ($N=400$, $\sigma_{u}=1$), but with $\phi=1.2$. Tables D.2 and D.3 report the results for the remaining combinations with $N=400$ and $\sigma_{u}=0.1$.
Tables D.4D.7 show the results of the simulations when the number of units in the panel reduces to $N=200$.
Table D.1: Simulation results for $N=400$, $\phi=1.2$ and $\sigma_{u}=1$
@@
Table D.2: Simulation results for $N=400$, $\phi=0.8$ and $\sigma_{u}=0.1$
@@
Table D.3: Simulation results for $N=400$, $\phi=1.2$ and $\sigma_{u}=0.1$
@@
Table D.4: Simulation results for $N=200$ units, $\phi=0.8$ and $\sigma_{u}=1$
@@
Table D.5: Simulation results for $N=200$ units, $\phi=0.8$ and $\sigma_{u}=0.1$
@@
Table D.6: Simulation results for $N=200$ units, $\phi=1.2$ and $\sigma_{u}=1$
@@
Table D.7: Simulation results for $N=200$ units, $\phi=1.2$ and $\sigma_{u}=0.1$
@@ Appendix E Additional application material
Table E.1: Definition of the variables included in the initial dataset
@@
Standardized math test score Standardized math test score of fifth-grade students. School year 2012/2013-school year 2020/2021 National Institute for the Evaluation of Educational Instruction and Training (INVALSI)
@@
Economic classification dummies Without specialization, non-manufacturing (touristic), non-manufacturing (non-touristic), made in Italy, other manufacturing 2011 Italian National Institute of Statistics (Istat)
Dummy district LLM with at least one industrial district 2011 Istat
Regional dummies A dummy for each of the 20 Italian regions Istat
Population dummies $\leq$ 10,000; (10,000; 50,000]; (50,000; 100,000]; (100,000; 500,000]; >500,000 2011 Istat
Share of individuals with a university degree Share of individuals aged 30-34 with a university degree 2014 Istat
Share of individuals with a high-school degree Share of individuals aged 30-34 with a high-school degree 2014 Istat
Share of urban area Urban surface / Total surface 2012 Istat
Share of population in the periphery Share of population located in municipalities considered as peripheral or ultra-peripheral according to the SNAI classification 2011 Istat
Number of public libraries Number of public libraries 2016 Istat
Number of visitors to museums Number of visitors to museums, galleries, archaeological sites, and monuments per 100,000 inhabitants 2015 Istat
Average maximum temperature 30-year average (1980-2010) of maximum air temperature (°C) 1980-2010 CAMS European Air Quality Reanalysis Data
Average minimum temperature 30-year average (1980-2010) of minimum air temperature (°C) 1980-2010 CAMS European Air Quality Reanalysis Data
Average total yearly precipitation 30-year average (1980-2010) of total precipitation (mm) 1980-2010 CAMS European Air Quality Reanalysis Data
Turnout referendum Turnout at the constitutional referendum that was held in Italy on 4 December 2016 2016 Ministry of the Interior
Share of Yes at the referendum Share of Yes at the constitutional referendum that was held in Italy on 4 December 2016 2016 Ministry of the Interior
Percentage of children aged 0-2 years who have utilized childcare services Children aged 0-2 years who have utilized childcare services provided by Municipalities (nurseries, micro-nurseries, or integrated and innovative services) / Resident children aged 0-2 years * 100. 2016 Istat
@@
Standardized Italian test score Standardized Italian test score of fifth-grade students. School year 2012/2013-school year 2018/2019 INVALSI
Percentage of participants to the math test Percentage of participation in the math test (pupils who took the test out of expected pupils) School year 2012/2013-school year 2018/2019 INVALSI
Percentage of participants to the Italian test Percentage of participation in the Italian test (pupils who took the test out of expected pupils) School year 2012/2013-school year 2018/2019 INVALSI
Average percentage score for math Average percentage score for math - adjusted for cheating School year 2012/2013-school year 2018/2019 INVALSI
Average percentage score for Italian Average percentage score for Italian - adjusted for cheating School year 2012/2013-school year 2018/2019 INVALSI
Standard deviation for math Standard deviation of the percentage score for math School year 2012/2013-school year 2018/2019 INVALSI
Standard deviation for Italian Standard deviation of the percentage score for Italian School year 2012/2013-school year 2018/2019 INVALSI
Log of the Gini index Log of the Gini index 2013-2019 Ministry of Economy and Finance (MEF)
Share of foreign population Foreigners / population 2013-2019 Istat
Unemployment rate Resident population aged 15+ not in employment but currently available for work / Labor force 2013-2019 Istat
Share of graduate mayors Share of municipalities with a mayor with a university degree 2013-2019 Ministry of the Interior
Share of recycled waste Share of recycled waste 2013-2019 Italian National Institute for Environmental Protection and Research (Ispra)
Share of workers in manufacturing Share of workers in manufacturing 2013-2019 Istat
Share of old population Share of population aged >=65 2013-2019 Istat
Declared income per capita Declared income per capita 2013-2019 MEF
Log of declared income (total) Log of declared income (total) 2013-2019 MEF
Share of income for pensions Share of overall declared income for pensions 2013-2019 MEF
Share of income for employment Share of overall declared income for employment 2013-2019 MEF
Population Resident population 2013-2019 Istat
Average price per square meter Average price per square meter - house 2013-2019 Osservatorio del Mercato Immobiliare – Agenzia delle Entrate
Notes. In our application, time-invariant covariates are either features of the LLM not subject to time changes (e.g. area of the LLM) or features taken at a time before 2017. Data from the 2020/2021 school year represent the first available post-pandemic data point for the standardized math test score, as there was no test for the school year 2019/2020 due to the national lockdown. Following Carlana et al. (2023), math test scores have been standardized for the entire dataset using the 2013-2019 values as a benchmark, with a mean of 0 and a standard deviation of 1.
Table E.2: Performances with different numbers of variables
@@
Table E.3: Definition of the 20 variables included in the estimation process (random forest)
@@
Note: These are the 20 variables selected via the panel CV procedure by the preliminary random forest.
Table E.4: Definition of the variables included in the data-driven CATEs analysis
@@
COVID-19 impact on the standardized math test score (dependent variable) Treatment effect of the COVID-19 crisis on the standardized math test score (SD) School year 2020/2021 Estimated via the MLCM with random forest
Share of income accruing to low-income earners Share of income accruing to individuals with an overall income $\leq$ €26,000 2019 MEF
Average download speed of the internet Average download speed of the internet 2018 AGCOM
Share of households without internet access Share of households without internet access 2018 AGCOM
Tourism LLM Dummy variable equal to 1 for LLMs specialized in tourism 2011 Istat
Share of individuals with a university degree Share of individuals aged 30-34 with a university degree 2015 Istat
Population Resident population 2019 Istat
Per capita expenditure on gambling Per capita expenditure on gambling 2019
@@
Unemployment rate Resident population aged 15+ not in employment but currently available for work 2019 Istat
@@
Excess mortality estimates Municipality-level excess mortality estimated by applying ML techniques to all-cause deaths data, aggregated at the LLM level From Feb 21, 2020 to Sep 30, 2020 Cerqua et al. (2021)
@@
Share of temporary contracts Number of employees with temporary contracts in October divided by the number of employees in October 2015 Istat
@@
Share of jobs in suspended economic activities Share of jobs in activities suspended in March 2020 by the Italian Government due to the spread of the pandemic 2017 Istat
Per capita income The amount of money earned per person 2019 MEF
@@
Share of firms having employees in short-time work subsidies The number of firms with employees receiving short-time work subsidies divided by the universe of firms registered in the Business Register Average (2015-2018) Ministry of Labor and Social Policies
@@
Share of population living in peripheral areas Share of population living in areas defined by Istat as peripheral or ultra-peripheral Jan 1, 2020 Istat
@@
Dependency ratio The ratio of those typically not in the labor force (the dependent part, ages 0 to 14 and 65+) and those typically in the labor force (the productive part, ages 15 to 64) Jan 1, 2020 Istat
Gini index Gini index 2019 MEF
@@
Average price per square meter Average price per square meter - house 2019 Osservatorio del Mercato Immobiliare – Agenzia delle Entrate
Figure E.1: Variable importance ranking – Preliminary random forest.
Figure E.2: Unit-level panel placebo test by year.
Figure E.3: SHAP values of the counterfactual 2020 forecasts produced by the final selected model (random forest) for a random 20% subsample of LLMs.
Appendix F Replication study
In this section, we revisit the empirical application of Callaway and Sant’Anna (2021) on the impact of minimum wage policy on teen employment in the United States (US). The main goal of this exercise is to assess whether the MLCM can replicate findings from a study that used a popular method based on the canonical use of control units. Additionally, we demonstrate how the MLCM can be easily adapted to a setting with staggered adoption of treatment. For the purpose of this replication, we simulate an environment in which the researcher has access to but cannot employ untreated units due to violations of the no-interference assumption. We check the reliability of these estimates by using the results of Callaway and Sant’Anna (2021), obtained via their difference-in-differences estimator for staggered adoption contexts with multiple time periods, as the ground truth.
We use county-level data for the period 2001-2007 from Callaway and Sant’Anna (2021) and apply only minimal pre-processing before proceeding with estimation. First, we discard never-treated units. Second, we employ the same set of variables included by Callaway and Sant’Anna (2021) in their specification with covariates: region dummies, county population, county median income, the fraction of white population that is white, the fraction of the population with a high school education, the county’s poverty rate. Third, we also include, as additional predictors, a lag of the outcome variable (the log of teen employment), county fixed effects—using the approach recommended by Johannemann et al. (2019)—and region-year dummies.
We then proceed to estimate the effects of staggered implementation of minimum wage policy on teen employment by forecasting counterfactuals with the MLCM. The key point to keep in mind is that not only we discard never-treated units, but we also do not rely on the available not-yet-treated units to build counterfactuals. Regarding the identification assumptions, we assume additivity and no anticipation, as in Callaway and Sant’Anna (2021), but we replace their parallel trends assumption with our Assumption 4 that implicitly rules out concomitant shocks and policies. This assumption appears reasonable given the context: the years under analysis are those just before the Great Recession, during which no other major employment shock occurred. [^31] Youth unemployment trends remained relatively stable in the US during these years, with a consistent value of 10.5% in both 2001 and 2007 (the initial and final years of the dataset); see data from the World Bank here. We also replace their standard SUTVA assumption, which postulates no interference among units, with our weaker version of SUTVA (Assumption 1), which only requires no hidden versions of the treatment.
We employ two versions of the MLCM: one with a fully non-linear model, random forest, and the other with a linear model, LASSO. LASSO outperformed random forest on this specific dataset across all treatment cohorts, making it the chosen method for forecasting counterfactual teen employment levels in the post-treatment periods. [^32] For this reason, we do not need to rely on Assumption 5, which is only required for identification in non-linear multi-step ahead forecasting.
The results of this no-control replication exercise are reported in the event-study graph of Figure F.1 below. This graph illustrates group-time average treatment effects (i.e., the estimand proposed by Callaway and Sant’Anna (2021)) and should be compared with the two sets of estimates reported in Figure 1 of the original study.
Figure F.1: Group-time average treatment effects – Replication without controls of Callaway and Sant’Anna (2021). Note: Confidence intervals estimated using 1000 block-bootstrap replications. To ease comparability, the range of the Y axis is the same used by Callaway and Sant’Anna (2021).
Despite completely discarding information from untreated units, we closely replicate their findings: the sign of the post-treatment estimates is always consistent (with the sole exception of the 2006 effect for the cohort first treated in 2006), while the magnitude differs to some extent but consistently falls within the range of confidence intervals from the two specifications of Callaway and Sant’Anna (2021) with conditional and unconditional parallel trends. The qualitative story remains unchanged, pointing to negative impacts of the minimum wage increase on county-level teen employment: our estimate of the ‘global’ average treatment effect is –3% (95% confidence intervals: –3.8; –2), which aligns with the estimate reported by Callaway and Sant’Anna (2021) of –3.2% for their specification with conditional parallel trends. [^33] Callaway and Sant’Anna (2021) refer to this estimate as the ‘simple weighted average’. It is the weighted average (by group size) of all available group-time average treatment effects. Lastly, pre-treatment estimates are all indistinguishable from zero. This is an improvement over the original study, where some of the placebo estimates were relatively large and statistically significant, casting doubts on the validity of the parallel trends assumption (Callaway and Sant’Anna, 2021).
Moreover, our county-specific treatment effects can be exploited to investigate patterns of treatment effect heterogeneity. One approach is to leverage these unit-level estimates by mapping them to explore potential spatial patterns. Figure F.2 shows county-specific effects, revealing highly heterogeneous estimates across states. Notably, the minimum wage had a particularly strong negative impact in Michigan, which is also the state that increased the minimum wage the most during the period under analysis (Callaway and Sant’Anna, 2021).
Figure F.2: County-specific treatment effects of the minimum wage on teen employment. Notes: These estimates are individual treatment effects averaged over the post-treatment period. The final sample used by Callaway and Sant’Anna (2021) includes county-level teen employment data for 29 states, matched with county characteristics. Of the 29 states considered in the Callaway and Sant’Anna (2021) analysis, 13 are treated. These 13 states comprise 930 counties. However, due to missing data in the original dataset, the number of treated counties is 907.
This replication study demonstrates the flexibility and applicability of the MLCM in staggered adoption settings or in cases with potential spillovers and general equilibrium effects contaminating the outcomes of untreated units. It also highlights that the MLCM can be used as a robustness check for traditional identification strategies, offering an alternative estimation method to assess the plausibility of the key identification assumptions, validate the main estimates, and rule out spillover effects. Finally, the MLCM generates individual treatment effects, providing a more granular building block compared to the group-time average effects of Callaway and Sant’Anna (2021) and most other traditional estimators leveraging control units, thus allowing for a more comprehensive investigation into the heterogeneity of impacts.